OpenStudy (anonymous):

Solve the following equation. Be sure to check the answer in the original equation if you multiply both sides by an expression that contains the variable.

OpenStudy (anonymous):

\[\frac{ 3 }{ x-6 } - \frac{ 2x }{ x+1 }=\frac{ -54 }{ x^2-5x-6 }\]

OpenStudy (anonymous):

So in this instance, I would multiply both sides by the LCD of the fractions.

OpenStudy (anonymous):

Don't I have to factor out the last denominator?

OpenStudy (anonymous):

Not necessarily. You can if you want to, but it's not strictly necessary. You'll see when you find the LCD.

OpenStudy (anonymous):

Ok. I'll do it your way cuz math is too frustrating right now.

OpenStudy (anonymous):

Have you determined the LCD?

OpenStudy (anonymous):

uhhh, (x-6) (x+1)?

OpenStudy (anonymous):

right. so what happens to the denominator on the right side?

OpenStudy (anonymous):

you multiply it by x-6?

OpenStudy (anonymous):

well you're multiplying everything by (x-6) (x+1)

OpenStudy (anonymous):

so since the denominator is (x-6) (x+1), it goes away

OpenStudy (anonymous):

So, I'm left with -54

OpenStudy (anonymous):

Do all the (x-6)(x+1)s cancel out or just the last ones?

OpenStudy (anonymous):

I'm closing the question just to post another question while I get help with this one...

OpenStudy (anonymous):

@sourwing , can you help us?

OpenStudy (anonymous):

Yeah. sorry for being afk. On the right side you're left with -54. What about on the left?

OpenStudy (anonymous):

-2x^2-15x-3?

OpenStudy (anonymous):

hm.... check your signs...

OpenStudy (anonymous):

your numbers are right though

OpenStudy (anonymous):

- & + confuse me sometimes

OpenStudy (anonymous):

lol yeah just keep it clear in your head >.>

OpenStudy (anonymous):

I get confused sometimes too...

OpenStudy (anonymous):

lol good cuz I'm not the only one lol

OpenStudy (anonymous):

right so write it out 3(x+1)-(2x)(x-6) (3x+3)-(2x^2-12) 3x+3-2x^2+12 -2x^2+3x+15

OpenStudy (anonymous):

set that equal to -54. Do you know how to get the answer from there?

OpenStudy (anonymous):

yes! subtract 15 from both sides!

OpenStudy (anonymous):

um...actually, add 54 to both sides.

OpenStudy (anonymous):

wouldn't it be the same thing?

OpenStudy (anonymous):

oh never mind lol

OpenStudy (anonymous):

not necessarily. you want to set your expression equal to zero if you want to factor it and find the roots. if you add 15 to both sides, you would not find the roots of the equation.

OpenStudy (anonymous):

2x^2+3x+69=0

OpenStudy (anonymous):

yup. What are the roots?

OpenStudy (anonymous):

I kind of forgot how to find those >.< lol

OpenStudy (anonymous):

actually, I'm wrong. It's 2x^2+3x+57.

OpenStudy (anonymous):

There's no way that this thing has rational roots. You're going to have to use the quadratic formula. >.<

OpenStudy (anonymous):

I can use: \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Right?

OpenStudy (anonymous):

yeah. It's a quadratic equation, after all.

OpenStudy (anonymous):

that can't be right ...so far, I have \[x=\frac{ -3\pm \sqrt{-447} }{ 4 }\] Square roots cant be negative, right?

OpenStudy (anonymous):

um...it's -2x^2. Sorry x.x

OpenStudy (anonymous):

lol omg, ok

OpenStudy (anonymous):

ok, so now i have \[x=\frac{ -3\pm \sqrt{465} }{ -4 }\] Right?

OpenStudy (anonymous):

b is 15, a is -2, and c is 57

OpenStudy (anonymous):

try with those numbers

OpenStudy (anonymous):

I still end up with a negative square root

OpenStudy (anonymous):

hmm... you shouldn't. 225-(-8)(57) isn't less than one...

OpenStudy (anonymous):

oooh ok, I got 681

OpenStudy (anonymous):

yup. that seems all right.

OpenStudy (anonymous):

you should get two roots. what are they?

OpenStudy (anonymous):

I'm horrible at this part lol especially with big numbers

OpenStudy (anonymous):

that's fine. I've figured out the numbers.

OpenStudy (anonymous):

lol good, what are they?

OpenStudy (anonymous):

\[x=\frac{ 15\pm \sqrt{681} }{ 4 }\]

OpenStudy (anonymous):

It's good to be able to know the steps, but to check your answer, I'd recommend www.wolframalpha.com

OpenStudy (anonymous):

I tried that, I don't know what the hell it was trying to give me lol

OpenStudy (anonymous):

lol. well try those roots in the equation, and that's your work.

OpenStudy (anonymous):

I probably put it in wrong lol

OpenStudy (anonymous):

that site is really picky about parentheses

OpenStudy (anonymous):

yessssss lol

OpenStudy (anonymous):

lol does that solve your problem?

OpenStudy (anonymous):

I need to go soon, so if you have any questions now is a good time.

OpenStudy (anonymous):

It says I need 2 answers

OpenStudy (anonymous):

you do have 2 answers. it's plus/minus. one answer is 15+√681 and one is 15-√681 (the numerators that is)

OpenStudy (anonymous):

Nevermind, I got it!!

OpenStudy (anonymous):

all right. Bye!

OpenStudy (anonymous):

Thank you! Bye!